How do you find the GCF and LCM for 90xy^2z^3, 300x^2y^2z^2, 120z^4?
Answer by stanbon(75887)

How do you find the GCF and LCM for 90xy^2z^3, 300x^2y^2z^2, 120z^4?
------------------
Prime Factor Forms:
90 = 2*3^2*5
300= 2^2*3*5^2
120= 2^3*3*5
------------------
GCF: 2*3*5 =30
LCM: 2^3*3^2*5^2 = 1800
---------------------------
Final Answer:
GCF = 30z^2
LCM = 1800x^2y^2z^4
========================
Cheers,
Stan H.